Exercise 14-8 Preparing a schedule of cash payments for inventory purchases LO 14-3 s require Stuart . that 90 of a 4 of 6
Answer #1

a)

Schedule of Cash Payments for Inventory Purchases:

April May June
Payment for current accounts payable $96,300 ($127,000 * 90% = $114,300 ($139,000 * 90%) = $125,100)
Payment for previous accounts payable 10,000 ($107,000 * 10% = $10,700) $127,000 * 10% = $12,700)
Total budgeted payments for inventory $106,300 ($114,300 + $10,700 = $125,000) ($125,100 + $12,700) = $137,800)

b)

Accounts payable the company will report on its pro forma balance sheet at the end of the second quarter will be the June purchases ($139,000) * 10% = $13,900.